4^√8 x 4^√16
--(divide)--
8^√2 x 8^√3

Answer by Imgone(6)

I'm going to be rounding the square roots to the nearest hundredths place. Square root of 8 is 2.83. So 4^2.83 = 50.56. Square root of 16 is 4 so 4^4 = 256. So 50.56 multiplied by 256 = 12943.36. Square root of 2 is 1.41. Square root of 3 is 1.73. So 8^1.41 multiplied by 8^1.73 = 685.105.
12943.36 divided by 685.105 = 18.89252012465243
Your final answer is 18.89252012465243 units.